A satellite circles the earth in an orbit whose ra
Answer #1

The velocity of the rotation v = sqrt (G M / r )
Where G is the gravitational constant ,M is the mass of the earth , r is the radius of the orbit
r = 4.67 x R = 4.67 x 6.38 x106 = 29.7946 x 106 m
v = sqrt (6.67 x 10-11 N m2 / kg2 x 5.98 x 1024 kg /   29.7946 x 106 m)
v = 3.659 x103 m /s
The angular velocity is
\omega = v / r
\omega  = 3.659 x 103 / 29.7946 x 106
\omega  = 0.1228 x 10-3 s-1
T = 2 \pi / \omega   
T = 2 x 3.14 /  0.1228 x 10-3 s-1
T = 51140.06 s